A blue ball is thrown upward with an initial speed of 20 m/s, from a height of 0.8 meters above the ground. 2.4 seconds after the blue ball is thrown, a red ball is thrown down with an initial speed of 5.3 m/s from a height of 22.4 meters above the ground. The force of gravity due to the earth results in the balls each having a constant downward acceleration of 9.81 m/s2. 1) How long after the blue ball is thrown are the two balls in the air at the same height?
Answer #1

Given the height:

y = y_{0} + v_{0}t - \frac{1}{2}gt^{2}

Since the heoght is the same, they balls can be related as:

y_{0_{B}} + v_{0_{B}}t - \frac{1}{2}gt^{2} = y_{0_{R}} + v_{0_{R}}(t - 2.4)- \frac{1}{2}g(t-2.4)^{2}

y_{0_{B}} + v_{0_{B}}t - \frac{1}{2}gt^{2} = y_{0_{R}} + v_{0_{R}}(t - 2.4)- \frac{1}{2}g(t^{2}-4.8t+5.76)

v_{0_{B}}t - \frac{1}{2}gt^{2} - v_{0_{R}}t + \frac{1}{2}g(t^{2}-4.8t)= y_{0_{R}} - y_{0_{B}} - 2.4v_{0_{R}}- \frac{1}{2}g(5.76)

(v_{0_{B}} - v_{0_{R}} - \frac{1}{2}g(4.8))t = y_{0_{R}} - y_{0_{B}} - 2.4v_{0_{R}}- \frac{1}{2}g(5.76)

\small (20m/s - 5.3m/s - \frac{1}{2}(9.81m/s^{2})(4.8))t = 22.4m - 0.8m - 2.4(5.3m/s)- \frac{1}{2}(9.81m/s^{2})(5.76)

\small (-8.844)t = -19.3728

\small t =\frac{-19.3728}{(-8.844)}

ANSWER \small t = 2.191s

I hope it helps!!